OpenStudy (anonymous):

1/2x-1/3y+z=1 x-y+z=-2 x+2/3z=0 i only need help finding y, thanks!!!

jhonyy9 (jhonyy9):

so than there are x/2 -y/3 +z =1 x-y+z=-2 x+2z/3=0 so here have not y ?

OpenStudy (anonymous):

wouldnt equation c be plugged into equations and b

OpenStudy (anonymous):

* a and b

jhonyy9 (jhonyy9):

but real in 3rd have not y ?

jhonyy9 (jhonyy9):

or you have missed it ?

OpenStudy (anonymous):

no, i thought you would change it to x=2/3x+0

OpenStudy (anonymous):

1/2(2/3x+0)-1/3y+z=1

OpenStudy (anonymous):

thts equation a

jhonyy9 (jhonyy9):

please first of all rewrite me please your exercise correct ok ?

OpenStudy (anonymous):

it is correct

jhonyy9 (jhonyy9):

so than in the 3rd equation have not y yes ?

OpenStudy (anonymous):

it doesnt have y

jhonyy9 (jhonyy9):

ok

jhonyy9 (jhonyy9):

x/2 -y/3 +z=1 so what will be here the common denominator for you can assum these fractions ?

OpenStudy (anonymous):

yes,

jhonyy9 (jhonyy9):

can you assuming these fractions ?

jhonyy9 (jhonyy9):

wait

OpenStudy (anonymous):

 then...?